2. Calculate the equilibrium constant for the reaction

N¬2O4(g) 2NO2(g)


given that fGo (N2O4(g)) = 97.89 kJ mol-1 and fGo (NO2(g)) = 51.31 kJ mol-1

Calculate the equilibrium constant at 100oC given that fHo (N2O4(g)) = 9.16 kJ mol-1 and fHo (NO2(g)) = 33.18 kJ mol-1.
1
Expert's answer
2020-05-10T09:20:02-0400

N2O4(g) = 2 NO2(g)

Equilibrium constant K is:

K = exp(-ΔrG0/RT)

where ΔrG0 = 2*ΔfG0(NO2) - ΔfG0(N2O4) = 2*51.31 - 97.89 = 4.73 kJ,

T = 100 oC + 273 = 373 K

K = exp(-4730/(8.31*373)) = 0.217

Given the enthalpy values, the change of entropy can be calculated:

ΔrG0 = ΔrH0 - T*ΔrS0 = 2*ΔfH0(NO2) - ΔfH0(N2O4) - T*ΔrS0

4730 = 2*33180 - 9160 - 373*ΔrS0

ΔrS0 = 140.67 J/K

Answer: equilibrium constant K = 0.217